# VR-GS: A Physical Dynamics-Aware Interactive Gaussian Splatting System in Virtual Reality

### v1

@@ Abstract.
* indicates equal contributions.
Affiliations: ${}^{1}$, ${}^{2}$, ${}^{3}$tah, ${}^{4}$, ${}^{5}$tyle3D Research, ${}^{6}$, ${}^{7}$mazon
As consumer Virtual Reality (VR) and Mixed Reality (MR) technologies gain momentum, there’s a growing focus on the development of engagements with 3D virtual content. Unfortunately, traditional techniques for content creation, editing, and interaction within these virtual spaces are fraught with difficulties. They tend to be not only engineering-intensive but also require extensive expertise, which adds to the frustration and inefficiency in virtual object manipulation. Our proposed VR-GS system represents a leap forward in human-centered 3D content interaction, offering a seamless and intuitive user experience. By developing a physical dynamics-aware interactive Gaussian Splatting (GS) in a VR setting, and constructing a highly efficient two-level embedding strategy alongside deformable body simulations, VR-GS ensures real-time execution with highly realistic dynamic responses. The components of our VR system are designed for high efficiency and effectiveness, starting from detailed scene reconstruction and object segmentation, advancing through multi-view image in-painting, and extending to interactive physics-based editing. The system also incorporates real-time deformation embedding and dynamic shadow casting, ensuring a comprehensive and engaging virtual experience. Our project page is available at: https://yingjiang96.github.io/VR-GS/.
copyright: none
@@ 1. Introduction
The burgeoning growth in commercial VR and MR applications is propelling the quest for high-fidelity 3D content within the Computer Graphics (CG) and Computer Vision (CV) communities. Despite their capabilities of being rendered efficiently in VR, traditional 3D/4D content creation, reliant on 3D modeling tools and game engines, is time-intensive and complex, often beyond the reach of non-expert users. This accessibility barrier limits broader user engagement in high-quality content creation.
@@
Recognizing the advantages and limitations of the traditional graphics pipeline, our goal is to find a modern variant. To enhance the visual quality and ease of creation for non-expert users, we move away from traditional 3D models in graphics pipelines and instead, adopt state-of-the-art radiance field techniques for rendering. In this context, Neural Radiance Fields (NeRF) emerge as a natural choice. Despite its versatility, NeRF’s volume rendering falls short in efficiency for interactive rendering on Extended Reality (XR) devices, which demand high frame rates. Additionally, NeRF’s approach to handling deformations—requiring bending of query rays via an inverse deformation map—is slow (Pumarola et al., 2021). Fortunately, 3D Gaussian Splatting (GS) (Kerbl et al., 2023) has been recently introduced as an efficient and explicit alternative to NeRF. This method not only excels in rendering efficiency but also provides an explicit geometric representation that can be directly deformed or edited. The explicit nature of 3D GS can simplify the direct manipulation of geometry by solving a Partial Differentiable Equation (PDE) that governs the motions of Gaussian kernels. Furthermore, GS eliminates the need for high-fidelity meshes, UV maps, and textures, offering the potential for naturally photorealistic appearances. It is worth noting that many studies have already demonstrated the use of 3D GS for 4D dynamics (Park et al., 2021; Pumarola et al., 2021; Yang et al., 2023) and the integration of physics-based simulations for more realistic 4D content generation (Xie et al., 2023; Feng et al., 2023), as well as for creating animatable avatars (Zielonka et al., 2023; Xu et al., 2023b).
@@
In this paper, we introduce VR-GS, a physics-aware, interactive Virtual Reality (VR) system for immersive manipulation of 3D content represented with Gaussian Splatting (GS). To ensure an interactive experience, we utilize eXtended Position-based Dynamics (XPBD) (Macklin et al., 2016), a highly adaptable and unified physical simulator, for real-time deformation simulation. Direct incorporation of a simulator onto GS kernels presents challenges, as the simulation and rendering processes have distinct geometrical representations. To address this, we construct a tetrahedral cage for each segmented GS kernel group and embed these kernel groups into corresponding meshes. The deformed mesh, driven by XPBD, subsequently guides the deformation of the GS kernels. Noticing that simplistic embedding techniques can lead to undesirable, spiky deformations in GS kernels, we propose a novel two-level embedding approach. This method allows each Gaussian kernel to adapt to a smoothed average deformation of the surrounding tetrahedra. The intricate combination of GS and XPBD through our two-level embedding not only achieves real-time physics-based dynamics but also upholds high-quality, realistic rendering, enabling high-fidelity immersive VR experiences. In summary, our contributions include:
  • Immersive VR System Development: The development of a VR system blending physical dynamics with high-fidelity appearance for enriched user experiences. We extensively evaluate our framework with various experiments including user studies.
  • Real-Time 3D Content Interaction: The engineering of the system with a focus on human-centric, creative interaction with captured and virtual 3D content.
  • Comprehensive System Integration: The comprehensive combination of 3D GS, segmentation, image inpainting, and real-time physics-based solvers with a novel rendering geometry embedding algorithm.
@@ 2.1. Radiance Fields Rendering
A variety of 3D representations, such as mesh (Sorkine and Cohen-Or, 2004), point clouds (Qi et al., 2017), signed distance fields (Wang et al., 2021), and grids (Zhu and Bridson, 2005) were exploited in early work for visual computing tasks, including synthesis, estimation, manipulation, animation, reconstruction, and transmission of data about objects and scenes (Kerbl et al., 2023; Xie et al., 2022; Gao et al., 2023). Since neural radiance fields (NeRF) (Mildenhall et al., 2021) was proposed for novel view synthesis with differentiable volume rendering, it has been applied to versatile computer graphics and computer vision tasks, such as scene reconstruction (Oechsle et al., 2021; Meng et al., 2023; Yariv et al., 2021), synthesis (Sun et al., 2023; Huang et al., 2023a), rendering (Lombardi et al., 2021; Wu et al., 2023) and animation (Chen et al., 2021; Peng et al., 2021), simulation (Li et al., 2023a), etc. While Neural Radiance Fields (NeRF) have achieved remarkable image quality, they suffer from significant time and memory inefficiencies (Lindell et al., 2021). To mitigate these issues, various methods have been introduced. Sparse representations (Liu et al., 2020), decomposed strategies (Niemeyer and Geiger, 2021; Rebain et al., 2021), and multi-resolution encoding (Müller et al., 2022) have all been proposed to enhance volume rendering efficiency while maintaining high quality. However, NeRFs are implicit representations, making it challenging to detect and resolve collisions (Qiao et al., 2023). Recently, 3D Gaussian splatting (GS) proposed by Kerbl et al. (2023) utilizes an array of 3D Gaussian kernels to represent scenes explicitly, facilitating faster optimization and achieving state-of-the-art results.
@@ 2.2. Editing in Radiance-based Scene
A natural extension of NeRF is to support user-guided editing. Li and Pan (2023) made use of two proxy cages to offer interactive control of the shape deformation of NeRF. Besides geometry editing, text prompts (Wang et al., 2023) or a user-edited image (Bao et al., 2023) can also be adopted to conduct style transfer of NeRF. Besides, Lin et al. (2023) put forward exploiting 2D sketches to edit high-quality facial NeRF. Concerning NeRF texture editing, Huang et al. (2023a) utilized a coarse-fine disentanglement representation and a patch-matching algorithm to synthesize textures of different shapes from multi-view images. What’s more, De-Nerf (Wu et al., 2023) exploits a hybrid light representation that enables users to relight NeRF. In contrast, Gaussian Splatting is more appropriate for post-editing tasks thanks to its explicit nature and inspired a series of follow-up works (Chen et al., 2023; Fang et al., 2023; Ye et al., 2023; Duisterhof et al., 2023; Huang et al., 2023b).
@@
Another significant direction in NeRF research is the incorporation of dynamic details into static scenes. To achieve this, a popular strategy (Pumarola et al., 2021; Park et al., 2021) is to consider time as an additional input to the system, This method typically splits a dynamic NeRF into a canonical static field and a deformation field, with the latter mapping this canonical representation to a deformed one. More recently, Yang et al. (2023) drew inspiration from 3D GS to reconstruct a dynamic scene using 4D Gaussian primitives that change over time. However, the dynamics in these methods are generally confined to motions captured from input data, limiting their ability to synthesize unseen dynamics. To generate novel dynamics, Xie et al. (2023) and Feng et al. (2023) have integrated physics-based simulations into 3D static representations, paving the way for generating physically plausible and novel dynamic scenes.
2.3. Neural Radiance Fields in VR
@@
Rendering NeRF is computationally expensive for real-time VR applications, which causes high latency and low quality (Song et al., 2023; Li et al., 2022b). To address these challenges and enable high-fidelity rendering with minimal latency on a single GPU, various techniques have been proposed. These include gaze-contingent 3D neural representations (Deng et al., 2022), variable rate shading (Rolff et al., 2023), and hybrid surface-volume representations (Turki et al., 2023), all aimed at accelerating NeRF. In contrast, VR-NeRF (Xu et al., 2023a) leverages multiple GPUs to achieve high-quality volumetric rendering of NeRF. Beyond software acceleration techniques, RT-NeRF (Li et al., 2022a) utilizes an algorithm-hardware co-design framework to provide real-time NeRF solutions for immersive VR rendering. While these methods primarily focus on enhancing NeRF rendering, our work is dedicated to facilitating interactive editing within a VR environment.
@@
Li et al. (2023c) integrated affine transformation with NeRF to enable interactive features, such as exocentric manipulation, editing, and VR tunneling effects. However, their approach is limited to filter-based edits such as edge blurring and fails to support the deformation of virtual objects’ geometry. RealityGit (Li et al., 2023b) and Magic NeRF Lens (Li et al., 2023d) modify a NeRF model by erasing or revealing a portion of NeRF. Nevertheless, neither system possesses the functionality to deform the geometrical structure of the model as well. In contrast to these works which concentrate on interactive editing of NeRF, our paper introduces VR-GS, a new physical dynamics-aware interactive system for GS in VR, first enabling users to interact with 3D GS at a real-time rate.
3. Design Goals and Design Decisions
@@
We propose the development of a unified, physics dynamics-aware interactive VR system, meticulously designed for real-time interaction with Gaussian Splatting (GS).
@@
Figure 2. VR-GS is an interactive system meticulously crafted to integrate 3D Gaussian Splatting (GS) and eXtended Position-based Dynamics (XPBD) for generating a real-time interactive experience in Virtual Reality (VR). Beginning with multi-view images, the pipeline skillfully combines scene reconstruction, segmentation, and inpainting using Gaussian kernels. These kernels form the foundation for VR-GS’s utilization of the sparse volumetric data structure VDB, facilitating bounding mesh reconstruction and subsequent tetrahedralization. VR-GS further harnesses a novel two-level rendering geometry embedding, XPBD, collision detection, and shadow casting techniques, all converging to deliver a captivating and immersive user experience.
3.1. Design Goals
G1. Immersive and Realistic
@@ 3.1. Design Goals
A VR system should provide a deeply immersive 3D environment that convincingly simulates physical presence, where objects possess a high degree of realism (Kalawsky, 1999). Accordingly, our targeted VR system aims to offer users an experience that is both immersive and realistic, characterized by dynamic motions, 3D virtual shapes, and illuminations that closely mirror the real world. This goal sets our system apart, especially in how it handles physics-based 3D deformations and dynamics, as opposed to those that rely solely on geometric deformation.
G2. Real-Time Interaction
For a VR system, low latency is crucial to prevent disorientation or motion sickness (Sutcliffe et al., 2019; Deng et al., 2022). Slow system responses can severely impair the quality of interaction. Our proposed VR system is engineered to provide users with instantaneous output and feedback during their interaction. This includes real-time responsiveness in various operations such as transformation, rotation, scaling, duplication, undoing, as well as physical and geometric editing. To our knowledge, VR-GS represents the first interactive physics-based GS editor.
G3. Unified Framework
VR-GS exemplifies a holistic framework, skillfully integrating both rendering and simulation within a single unified system. Our design philosophy focuses on harmonizing various 3D representation formats, utilizing a blend of thoughtful algorithms and design principles. This integrated approach guarantees a seamless user experience, adhering to the concept of “what you see is what you simulate”(Müller et al., 2016).
G4. User-Centric Design
Essential to a user-friendly VR system is the integration of a flexible model or simulation that supports intuitive and task-oriented user interaction (Stanney et al., 2003). VR-GS is designed to provide this intuitive control, along with responsive feedback, ensuring a natural and seamless interaction experience. The system’s operations and outputs are meticulously aligned with user expectations and commonsense understandings (Hale and Stanney, 2014; Gabbard et al., 1999). Furthermore, VR-GS’s scene in-painting capability, which effectively fills gaps resulting from object removal, is a testament to its commitment to meeting user expectations.
3.2. Design Decisions
Our design decisions align closely with the design goals.
D1. Physics-Based Generative Dynamics
The VR-GS system leverages physics-based simulation techniques to produce realistic dynamics and facilitate 3D shape editing. Unlike methods that reconstruct motion from time-dependent datasets or utilize generative AI to drive Gaussian models, our system adheres to physical principles.
D2. Real-Time Editing and Manipulation
@@ 3.2. Design Decisions
To achieve both real-time performance and physically realistic editing and manipulation, we implement a reduced representation model and parallel-friendly algorithms within our frame budget. While the per-Gaussian-based discretization as done in (Xie et al., 2023) offers detailed dynamics, its time cost is impractical for our system’s needs. Instead, we utilize a tetrahedral mesh embedding reconstructed from Gaussian kernels to drive GS motion. We adopt XPBD with finite-element constraints to achieve real-time simulation paradigm prevalent in gaming and VR applications.
D3. Unified Simulation and Rendering
@@
Our system employs a hybrid approach, combining cage mesh and Gaussian kernels, to achieve real-time performance. To enhance alignment between these two representations, we focus on integrating their functionalities. The deformation gradient field from the cage mesh is adept at embedding a volumeless point cloud. However, as each Gaussian kernel resembles a rotated ellipsoid, it can lead to spiky artifacts during large deformation, a challenge highlighted in (Xie et al., 2023). To mitigate this, we employ a two-level interpolation strategy: initially embedding each Gaussian within an individual tetrahedron, followed by embedding these tetrahedra within the cage mesh to elevate the smoothness of the deformation field.
D4. Natural User Control and Prepossession
Diverging from traditional 2D interface inputs, VR-GS harnesses 3D input mechanisms for interacting with 3D objects. Moreover, the system incorporates object segmentation and scene inpainting functionalities to seamlessly fill gaps created by object removal in the GS training process. This integration ensures a consistent and coherent user experience.

### v2

@@ Abstract.
As 3D content becomes increasingly prevalent, there’s a growing focus on the development of engagements with 3D virtual content. Unfortunately, traditional techniques for creating, editing, and interacting with this content are fraught with difficulties. They tend to be not only engineering-intensive but also require extensive expertise, which adds to the frustration and inefficiency in virtual object manipulation. Our proposed VR-GS system represents a leap forward in human-centered 3D content interaction, offering a seamless and intuitive user experience. By developing a physical dynamics-aware interactive Gaussian Splatting (GS) in a Virtual Reality (VR) setting, and constructing a highly efficient two-level embedding strategy alongside deformable body simulations, VR-GS ensures real-time execution with highly realistic dynamic responses. The components of our system are designed for high efficiency and effectiveness, starting from detailed scene reconstruction and object segmentation, advancing through multi-view image in-painting, and extending to interactive physics-based editing. The system also incorporates real-time deformation embedding and dynamic shadow casting, ensuring a comprehensive and engaging virtual experience.
Gaussian Splatting, Neural Radiance Fields, Real-Time Interactions
submissionid: 608copyright: nonejournalyear: 2024copyright: acmlicensedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24; July 27-August 1, 2024; Denver, CO, USAbooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ’24 (SIGGRAPH Conference Papers ’24), July 27-August 1, 2024, Denver, CO, USAdoi: 10.1145/3641519.3657448isbn: 979-8-4007-0525-0/24/07ccs: Computing methodologies Virtual reality
@@ 1. Introduction
As digital technology advances, the importance of 3D content is becoming increasingly prominent across various industries, from entertainment to education. This growth is driving the demand for high-fidelity 3D content within the Computer Graphics (CG) and Computer Vision (CV) communities. Despite their capabilities of being rendered efficiently, traditional 3D/4D content creation, reliant on 3D modeling tools and game engines, is time-intensive and complex, often beyond the reach of non-expert users. This accessibility barrier limits broader user engagement in high-quality content creation.
@@
Recognizing the advantages and limitations of the traditional graphics pipeline, our goal is to find a modern variant. To enhance the visual quality and ease of creation for non-expert users, we move away from traditional 3D models in graphics pipelines and instead, adopt state-of-the-art radiance field techniques for rendering. In this context, Neural Radiance Fields (NeRF) emerge as a natural choice. Despite its versatility, NeRF’s volume rendering falls short in efficiency for interactive applications, which demand high frame rates. Additionally, NeRF’s approach to handling deformations—requiring bending of query rays via an inverse deformation map—is slow [Pumarola et al., 2021]. Fortunately, 3D Gaussian Splatting (GS) [Kerbl et al., 2023] has been recently introduced as an efficient and explicit alternative to NeRF. This method not only excels in rendering efficiency but also provides an explicit geometric representation that can be directly deformed or edited. The explicit nature of 3D GS can simplify the direct manipulation of geometry by solving a Partial Differentiable Equation (PDE) that governs the motions of Gaussian kernels. Furthermore, GS eliminates the need for high-fidelity meshes, UV maps, and textures, offering the potential for naturally photorealistic appearances. It is worth noting that many studies have already demonstrated the use of 3D GS for 4D dynamics [Park et al., 2021; Pumarola et al., 2021; Yang et al., 2023] and the integration of physics-based simulations for more realistic 4D content generation [Xie et al., 2024; Feng et al., 2024b], as well as for creating animatable avatars [Zielonka et al., 2023; Xu et al., 2023b].
@@
In this paper, we introduce a physics-aware interactive system for immersive manipulation of 3D content represented with GS. To ensure an interactive experience, we utilize eXtended Position-based Dynamics (XPBD) [Macklin et al., 2016], a highly adaptable and unified physical simulator, for real-time deformation simulation. Direct incorporation of a simulator onto GS kernels presents challenges, as the simulation and rendering processes have distinct geometrical representations. To address this, we construct a tetrahedral cage for each segmented GS kernel group and embed these kernel groups into corresponding meshes. The deformed mesh, driven by XPBD, subsequently guides the deformation of the GS kernels. Noticing that simplistic embedding techniques can lead to undesirable, spiky deformations in GS kernels, we propose a novel two-level embedding approach. This method allows each Gaussian kernel to adapt to a smoothed average deformation of the surrounding tetrahedra. The intricate combination of GS and XPBD through our two-level embedding not only achieves real-time physics-based dynamics but also upholds high-quality, realistic rendering. In summary, our contributions include:
  • A Physics-Aware Interactive System: Development of a system that enables interactive, physics-aware manipulation of 3D content represented with GS.
  • Two-Level Deformation Embedding: Introduction of a novel two-level embedding approach that allows Gaussians to adapt smoothly to the mesh, enhancing the deformation realism and preventing undesirable spiky artifacts.
We deploy our system on VR devices, enriched by segmentation, inpainting, and shadow map, offering users a rich platform for 3D content manipulation.
@@ 2.1. Radiance Fields Rendering
A variety of 3D representations, such as mesh [Sorkine and Cohen-Or, 2004], point clouds [Qi et al., 2017], signed distance fields [Wang et al., 2021], and grids [Zhu and Bridson, 2005] are exploited in early work for visual computing tasks, including synthesis, estimation, manipulation, animation, reconstruction, and transmission of data about objects and scenes [Kerbl et al., 2023; Xie et al., 2022; Gao et al., 2023]. Since neural radiance fields (NeRF) [Mildenhall et al., 2021] was proposed for novel view synthesis with differentiable volume rendering, it has been applied to versatile computer graphics and computer vision tasks, such as scene reconstruction [Oechsle et al., 2021; Meng et al., 2023; Yariv et al., 2021], synthesis [Sun et al., 2023; Huang et al., 2023a], rendering [Lombardi et al., 2021; Wu et al., 2023], interactive games [Xia et al., 2024] and animation [Chen et al., 2021; Peng et al., 2021], simulation [Li et al., 2023a], etc. While Neural Radiance Fields (NeRF) have achieved remarkable image quality, they suffer from significant time and memory inefficiencies [Lindell et al., 2021]. To mitigate these issues, various methods have been introduced. Sparse representations [Liu et al., 2020], decomposed strategies [Niemeyer and Geiger, 2021; Rebain et al., 2021], and multi-resolution encoding [Müller et al., 2022] have all been proposed to enhance volume rendering efficiency while maintaining high quality. However, NeRFs are implicit representations, making it challenging to detect and resolve collisions [Qiao et al., 2023]. Recently, 3D Gaussian splatting (GS) proposed by [Kerbl et al., 2023] utilizes an array of 3D Gaussian kernels to represent scenes explicitly, facilitating faster optimization and achieving state-of-the-art results.
@@ 2.2. Editing in Radiance-based Scene
A natural extension of NeRF is to support user-guided editing. Li and Pan [2023] made use of two proxy cages to offer interactive control of the shape deformation of NeRF. Besides geometry editing, text prompts [Wang et al., 2023] or a user-edited image [Bao et al., 2023] can also be adopted to conduct style transfer of NeRF. Besides, Lin et al. [2023] put forward exploiting 2D sketches to edit high-quality facial NeRF. Concerning NeRF texture editing, Huang et al. [2023a] utilized a coarse-fine disentanglement representation and a patch-matching algorithm to synthesize textures of different shapes from multi-view images. De-Nerf [Wu et al., 2023] exploited a hybrid light representation that enables users to relight NeRF. In contrast, Gaussian Splatting is more appropriate for post-editing tasks thanks to its explicit nature and has inspired a series of follow-up works [Chen et al., 2023; Fang et al., 2023; Ye et al., 2023; Duisterhof et al., 2023; Huang et al., 2023b].
@@
Another significant direction in NeRF research is the incorporation of dynamic components into static scenes. To achieve this, a popular strategy [Pumarola et al., 2021; Park et al., 2021] is to consider time as an additional input to the system, This method typically splits a dynamic NeRF into a canonical static field and a deformation field, with the latter mapping this canonical representation to a deformed one. More recently, Yang et al. [2023] drew inspiration from 3D GS to reconstruct a dynamic scene using 4D Gaussian primitives that change over time. However, the dynamics in these methods are generally confined to motions captured from input data, limiting their ability to synthesize unseen dynamics. To generate novel dynamics, Xie et al. [2024] and Feng et al. [2024b] have integrated physics-based simulations into 3D static representations, paving the way for generating physically plausible and novel dynamic scenes.
2.3. Real-time Neural Radiance Fields
@@
Rendering NeRF is computationally expensive for real-time applications, such as VR, which causes high latency and low quality [Song et al., 2023; Li et al., 2022b]. To address these challenges and enable high-fidelity rendering with minimal latency on a single GPU, various techniques have been proposed. These include gaze-contingent 3D neural representations [Deng et al., 2022], variable rate shading [Rolff et al., 2023], and hybrid surface-volume representations [Turki et al., 2023], all aimed at accelerating NeRF. In contrast to single-GPU solutions, VR-NeRF [Xu et al., 2023a] leveraged multiple GPUs to achieve high-quality volumetric rendering of NeRF. Beyond software acceleration techniques, RT-NeRF [Li et al., 2022a] utilized an algorithm-hardware co-design framework to provide real-time NeRF solutions for immersive VR rendering. While these methods primarily focus on enhancing NeRF rendering, our work is dedicated to facilitating interactive editing.
@@
Li et al. [2023c] integrated affine transformation with NeRF to enable interactive features, such as exocentric manipulation, editing, and VR tunneling effects. However, their approach is limited to filter-based edits such as edge blurring and fails to support the deformation of virtual objects’ geometry. RealityGit [Li et al., 2023b] and Magic NeRF Lens [Li et al., 2023d] modified a NeRF model by erasing or revealing a portion of NeRF. Nevertheless, neither system permits deforming the geometrical structure of the model. In contrast to these works which concentrate on interactive editing of NeRF, our system enables users to interact with 3D GS in a physically realistic way in real time.
3. System Design
@@
We propose a unified physics dynamics-aware interactive VR system for real-time interactions with Gaussian Splatting (GS).
@@
Figure 2. VR-GS is an interactive system designed to integrate 3D Gaussian Splatting (GS) and eXtended Position-based Dynamics (XPBD) for generating a real-time interactive experience. Beginning with multi-view images, the pipeline combines scene reconstruction, segmentation, and inpainting using Gaussian kernels. These kernels form the foundation for VR-GS’s utilization of the sparse volumetric data structure VDB, facilitating bounding mesh reconstruction and subsequent tetrahedralization. VR-GS further harnesses a novel two-level Gaussian embedding, XPBD, collision detection, and shadow casting techniques, all converging to deliver a captivating and immersive user experience.
Immersive and Realistic Generative Dynamics
@@
Our targeted system aims to offer users an immersive and realistic experience, characterized by dynamic motions, 3D virtual shapes, and illuminations that closely mirror the real world [Kalawsky, 1999]. This goal sets our system apart, especially in how it handles physics-based 3D deformations and dynamics, as opposed to those that rely solely on geometric deformation. The proposed system leverages physics-based simulation techniques to produce realistic dynamics and facilitate 3D shape editing. Unlike methods that reconstruct motion from time-dependent datasets or utilize generative AI to drive Gaussian models, our system adheres to physical principles. To create an immersive experience, we choose VR as the interaction platform.
Real-Time Interaction
@@
For an interactive system, low latency is crucial to prevent disorientation or motion sickness [Sutcliffe et al., 2019; Deng et al., 2022]. To offer instantaneous output and feedback, most interactive systems support basic transformations including rotation, scaling, and translation. Our system additionally offers users real-time physics-based interaction. To our knowledge, VR-GS represents the first interactive physics-based GS editor. To achieve both real-time performance and physically realistic editing, we implement a reduced representation model and parallel-friendly algorithms within our frame budget. While the per-Gaussian-based discretization as done in PhysGaussian [Xie et al., 2024] offers detailed dynamics, its time cost is impractical for our system’s needs. Instead, we utilize a tetrahedral mesh embedding reconstructed from Gaussian kernels to drive GS motion. We adopt XPBD with finite-element constraints to achieve real-time simulations.
Unified Framework
@@
VR-GS is a framework that integrates rendering and simulation within a unified pipeline. Our design principles adhere to the goal of “what you see is what you simulate” [Müller et al., 2016]. Unlike Xie et al. [2024], our system employs a hybrid approach, combining cage mesh and Gaussian kernels, to achieve real-time performance. The deformation gradient field of Gaussian kernels is embedded in the cage mesh through piecewise constant interpolation. With each Gaussian kernel resembling a rotated ellipsoid, this simple embedding strategy leads to spiky artifacts during large deformation, a challenge highlighted in [Xie et al., 2024]. To mitigate this, we employ a two-level interpolation scheme: initially embedding each Gaussian within an individual bounding tetrahedron, followed by embedding these tetrahedra within the simulated cage mesh to elevate the smoothness of the deformation field.
Additional Components
Our framework is designed to construct VR-compatible interactive settings derived from real-world captures. Consequently, it introduces complexities associated with facilitating intuitive user interactions and delivering real-time feedback that meets user expectations and common sense perceptions [Stanney et al., 2003; Hale and Stanney, 2014; Gabbard et al., 1999]. In particular, when users engage with objects within a scene, they anticipate natural movements and realistic appearances from both the selected objects and the surrounding environment. Moreover, it is necessary to address voids when an object is removed from its supporting base. To manage these challenges, our system additionally integrates functionalities for object segmentation and scene inpainting.